1. Without looking up the actual value, rank the following metals in order of increasing melting point.
indium, cesium
2. Calcium crystallizes in a face-centered cubic unit cell with an edge length of 558.84 pm. Calculate the density of the metal. answer in g/mL
3.What is the density of KCl at 25.00 °C if the edge length of its fcc unit cell is 628 pm? answer in g/cm3
1
Expert's answer
2020-02-04T06:35:01-0500


1. The melting and boiling points of typical metals, as a rule, increase with a transition from bottom to top along a group of a periodic system of chemical elements. Therefore, first comes cesium, then indium.


2. We calculate the density by the formula: p = m / V

Since Calcium crystallizes in a face-centered cubic, we find the volume by the formula:

V = a3 = 558.843 * pm = 17.4 * 10-23cm3

4 atoms in a crystalline bond

Find the mass: 4 Ca atoms * 40.078 g / 6.022 * 1023 atoms = 160.312g/6.022 * 10-23 = 26.62 * 10-23 g

p = m / V = ​​26.62 * 10-23g /17.4 * 10-23cm3= 1.529g/cm3 = 1.529g/ml


3.Similarly, according to formulas 2 points:

V = a3 = 6283*pm = 24.7 * 10-23cm3

M=4*39.0983+4*35.453/6.022*1023=49.519*10-23

p = m / V = ​​49.519* 10-23 g /24.7 * 10-23cm3= 2.00g/cm3